Time of last reading:
Jan. 20, 2019, 10 p.m.
Nov. 27, 2017, 12:45 p.m.
Issue Aug. 1, 2018 - Aug. 1, 2018
O3 Sensor replaced around 3:45pmIssue Feb. 12, 2018 - Feb. 22, 2018
It has been brought to our attention by Environmental Instruments Ltd (manufacturers of AQMesh) that due to a bug in a server systems update that NO2 and O3 data from all AQMesh pods were effected from 7am on 12th February until 9am on 22nd February 2018. No other channels are affected. We caution you against using NO2 and O3 data for this period until the corrections have been applied. We urge users to re-call this data from Tuesday 27th when all readings will have been re-processed. The Urban Observatory apologises for the inconvenience.Issue April 18, 2018 - April 18, 2018
New deployment Holystone roundabout, central reservation on A186 nearest property postcode is: Cloverfield, West Allotment, Newcastle upon Tyne, NE27 0BE Lamp post 1AE28 Google 55.029411, -1.521943Issue Oct. 24, 2018 - Oct. 25, 2018
All AQ Meshes did not report any data between Weds 24th Oct 06:00 – Thurs 25th Oct 11:15, reason not yet knownIssue Nov. 16, 2018 - Nov. 16, 2018
Approx 15:00 - 16:00 : Changed the sampling frequency for gases for all AQ Meshes to be standard across all pods : P1 (Gas Reading Frequency) - Ten secs (unchanged) P2 (Gas Averaging Frequency) - One Minute (was mostly 15 mins) P3 (Gas Transmission Frequency) - 30 Mins (was One Hour)Issue Nov. 22, 2018 - Nov. 23, 2018
22/11/18 17:00 AirMonitors carrying out server maintenance at approx 17:00 (UTC) on Thursday November 22nd 23/11/10:00 server back. Possible data loss during this periodIssue Dec. 6, 2018 - Dec. 6, 2018
06/12/2018 Major disruption to the 02 Cellphone/Data network today. This will affect many sensors that rely on that network (eg Emotes)Issue Nov. 29, 2018 - Dec. 12, 2018
Gas sensor type:NO2 Sensor failed at 21:06 on 29 Nov 2018. Failed sensor serial number:202100007 (Date code 107) Serial number of new sensor:240500559 (Date code 098) Notified 05/12/18 Received 11/12/18 Replaced 12/12/18 Please allow 2-3 days for new sensor to stabilise.Issue Dec. 17, 2018 - Dec. 31, 2019
Data looks erratic (high 1000s) - impending 03 sensor failure. Awaiting confirmation of new 03 sensor from AirMonitors.Issue Jan. 15, 2019 - Jan. 16, 2019
15/01/19 02:15 - 16/01/19 Data lost from all sensors due to quue failure. This data may eventually get back-filled.
Sensor Maintenance Records: